Circular No. 8648

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF URANUS Further to IAUC 8177, the IAU Working Group for Planetary System

Nomenclature has approved the following new designations and names of satellites of Uranus:

Uranus XXII Francisco = S/2001 U 3
Uranus XXIII Margaret = S/2003 U 3
Uranus XXIV Ferdinand = S/2001 U 2
Uranus XXV Perdita  = S/1986 U 10
Uranus XXVI Mab  = S/2003 U 1
Uranus XXVII Cupid  = S/2003 U 2

SUPERNOVAE 2005mf AND 2005mg Two additional supernovae have been discovered on unfiltered CCD im-

ages: 2005mf by the THCA Supernova Survey (cf. IAUC 8647) and 2005mg by J. Newton and T. Puckett (cf. IAUC 8647). Additional magnitudes for 2005mf in UGC 4798: Nov. 25, [18.5; Dec. 26.58 UT, 17.3. SN 2005mf is not visible on TNT images from Oct. 31, Nov. 13, and Dec. 13. Additional magnitudes by Puckett for 2005mg in UGC 155: Dec. 1, [19.5; 27.99, 15.2.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005mf Dec. 25.67 9^{h}08^{m}42.^{s}33 +44^{o}48′51′′.4 17.4 5′′.9 W, 13′′.3 N
2005mg Dec. 27.08 0 16 44.21 + 7 04 19.8 16.2 0′′.3 E, 14′′.7 S

COMETS C/2005 U7, C/2005 U8, C/2005 Y1 (SOHO) Additional near-sun comets have been found on SOHO website images

(cf. IAUC 8644). C/2005 U7 and C/2005 U8 were very faint and diffuse Kreutz sungrazers. Non-group comet C/2005 Y1 was around 8th mag (and very hard to spot) in C3 images, but it was well visible with a circular, diffuse head in C2 images, where it reached mag ∼ 7.1 with a very faint tail ∼ 5′ long at ∼ 5R_{⊙} on Dec. 17.125 UT; C/2005 Y1 was again approaching mag ∼ 8 when it left the C2 field-of-view.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 U7 Oct. 24.496 13^{h}49.^{m}5 −12^{o}51′ C2  QY 2005-Y02
C/2005 U8 30.038 14 12.1 −14 58 C2 HS 2005-Y02
C/2005 Y1 Dec. 16.975 17 35.2 −23 29 C2/3 SF 2005-Y07
2005 December 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green